# Physics-informed machine learning for combustion: A review

### v2

@@ 8.1.1 Limitations reinforced by combustion
To deal with these difficulties, the methods introduced in Section 2.3 should be appropriately used and improved. For example, the Fourier multiscale embedding can be used to enhance the multiscale representation 130, which has already been adopted in PIML for combustion 307. However, the introduction of hyperparameters is another minor issue, which can be resolved if the scale information is known. Moreover, there is no guarantee that this method will be effective on problems with a wider range of scales. In such cases, it is necessary to combine the domain decomposition methods, which are particularly important for large-domain and long-time problems. Domain decomposition can also facilitate parallel computation, thus accelerating the simulation. It is recommended to use knowledge-based domain decomposition if the approximate distribution characteristics of the solution are known in advance. Examples include combustors with reactant, mixing, reaction, and product regions. Adaptive domain decomposition can also be proposed for more complex problems. Unlike common domain decomposition methods, for combustion problems, each subdomain can use not only different networks but also different physical laws, which is similar to the adaptive combustion models 50. To deal with the multiphysics issue, imposing hard constraints on certain physical laws can always enhance the physical consistency and generalization ability. As for the ill-conditioning issue, some preconditioning strategies should be adopted, such as the normalization. For a deeper understanding, the ill-conditioning analysis should also be conducted for PIML solving combustion PDEs, which can be enlightened by the papers mentioned in Section 2.3.4. Although the reaction stiffness has been successfully solved in reaction ODE problems (see Section 4.1), it has not been thoroughly solved in reacting flow problems. As summarized in Table 5, when using the direct finite-rate chemistry in PIML, the most detailed reaction mechanism is only a two-step mechanism. Other studies either discard, simplify, tabulate, pre-learn, or online-learn the reaction part. Among these, only tabulating and pre-learning can represent the known detailed mechanism and may serve as feasible ways to tackle the stiffness issue in reacting flows, as long as they are differentiable. In fact, tabulating and pre-learning with NNs can be viewed as preconditioning strategies and can be categorized as building surrogate models of chemical kinetics, as mentioned in Section 4.2, not for CFD, but for PIML.
@@ 8.1.2 Discrepancies between physics, (physical) model, and data
As exemplified in Sections 4 and 5, to deal with these discrepancies between physics, model, and data, three strategies can be adopted: (1) Use reduced models, i.e., keep only the certain models and discard the uncertain parts. For example, the mass/continuity equation was adopted in most studies, as shown in Table 5, because it is certain, simple, and do not involve source terms that may need to be modeled for other equations. For the same reason, the energy and species equations are discarded in many studies. (2) Use simplified models. Examples include using RANS or LES models for turbulence and using simplified reaction mechanisms for complex reactions. Filtered PDEs have been proven to be effective for PIML to handle real-world data 207. (3) Learn the models, especially the turbulence and reaction closures, which have been illustrated in References 312, 109, 110, 313. These works demonstrate that the potential of PIML in inverse problems is not limited to field reconstruction and parameter inference, but can also assist in physical modeling. In addition to designing strategies to address these discrepancies, it is also necessary to have methods to measure these discrepancies, i.e., uncertainty quantification (UQ) 216, 47. The methods for quantifying data and model uncertainties mentioned in Section 2.3.4 can be applied to PIML for combustion.
@@ 8.1.3 Reproducibility
Reproducibility is the foundation of academic research, ensuring the reliability, transparency, and implementability of the adopted methods in the papers, thereby guaranteeing the credibility of their conclusions. Many fields face reproducibility issues, or the so-called reproducibility crisis 439, 440, and we believe that PIML for combustion is no exception. According to Reference 441, overoptimism exists in ML for fluid-related PDEs, which is caused by weak baselines and reporting biases. The former does not apply to most studies on PIML for combustion because their goal is not to exceed a baseline model. However, reporting bias is highly likely to exist, which includes selective reporting, outcome switching, and publication bias 441. In addition to the reporting bias, the reproducibility issues can also be caused by other pitfalls, including the sensitivity to hyperparameters, ambiguity in implementation details, and inaccessibility of data and codes. As demonstrated in Tables 4 and 5, most studies on PIML for combustion do not share codes, which is very different from the open source ecosystem of the AI community.
To deal with the reproducibility issues, McGreivy et al. 441 call for bottom-up cultural changes to minimize biased reporting and top-down structural reforms to reduce perverse incentives. We briefly outline some solutions at the researcher level: (1) Provide detailed model configurations, parameter settings, and evaluation metrics. (2) Improve data accessibility by either making data available or using publicly available data, such as those in Table 6. When establishing public datasets, it is preferable to follow the principles of FAIR (Findability, Accessibility, Interoperability, and Reusability) 442. When establishing public benchmarks, some suggestions can be found in Reference 354. (3) It is preferable to provide the codes, as we have done in References 115, 319 and Listing 32. When developing public software, some suggestions can be found in Reference 443. (4) Report not only positive results but also negative results. (5) Conduct replicate experiments and report the mean and variance of the results to demonstrate that the impact of randomness is negligible. (6) Conduct hyperparameter sensitivity analysis. For hyperparameters that are highly sensitive, sufficient explanation should be provided. Finally, we emphasize that the reproducibility issues can also be alleviated by the establishment of an open, active, and rigorous community for PIML for combustion, which involves the standardization of processes, tools, and definitions, as well as the open sourcing of data, codes, and tools. There is still a long way to go in this regard.
@@ 8.2 Discussions on implementation choices
Forward vs inverse problems. As shown in Tables 4 and 5, PIML is currently used more for inverse problems than for forward problems, and this is not limited to the field of combustion. There are many reasons for this: (1) Inverse problems are usually simpler than forward problems when both types of problems can be studied because of the involvement of data. When solving forward problems fails, inverse problems are the only option. (2) Inverse problems are more common than forward problems in practice. Forward problems have strict requirements and need complete equations and IC/BCs, while inverse problems have more relaxed requirements. Therefore, it is not that researchers do not study forward problems, but that the specific problems themselves are inverse problems. (3) Inverse problems have richer application significance than forward problems in terms of field reconstruction, parameter inference, physical modeling, and kinetics discovery. Forward problems often only involve simulation, which can also be achieved by traditional numerical methods. (4) Inverse problems are better suited to leveraging the advantages of PIML than forward problems. It is well known that the overall performance of PIML in forward problems is far inferior to that of traditional numerical methods 441. However, one advantage of PIML is that it can easily couple data and incomplete equations to solve inverse problems, which is not straightforward for traditional numerical methods. In fact, the data-driven characteristic is an inherent advantage of ML, which can be leveraged in PIML. Nevertheless, the forward problems still have two important aspects: (1) In terms of academic research, the forward problem is the basis of the inverse problem. Research on the forward problems helps deepen our understanding of PIML, which in turn helps us understand the inverse problems. Furthermore, many proposed methods on the forward problems in terms of representation models, optimization processes, and PDE systems can be seamlessly transferred to the inverse problems. (2) In terms of practical applications, although PIML is inferior to traditional solvers in solving single PDE, it has demonstrated advantages in parametric/surrogate solving. Examples can be found in Section 2.3.4, Section 2.4, and Reference 307. Compared with traditional solvers, it is more efficient, provides a continuous representation in the parameter space, and can easily perform parameter sensitivity analysis. Note that parametric/surrogate modeling can also be implemented in the way of inverse problems, where the data can be utilized. To sum up, both types of problems are important. For researchers and engineers, the choice between forward problems and inverse problems should depend on the specific circumstances and objectives.
@@
Continuous vs discrete PINNs. As defined in Section 2.3.1, continuous PINNs follow the coordinate-to-variable mapping way, i.e., $\widehat{\mathbf{u}}=\mathrm{NN}(\mathbf{z})$, while discrete PINNs follow the field-to-field mapping way, i.e., $\widehat{\mathbf{U}}_{\mathrm{out}}=\mathrm{NN}(\mathbf{U}_{\mathrm{in}})$. Both have their own advantages and disadvantages 139. Continuous PINNs provide a continuous representation of the solution, making them mesh-independent and allowing for easy query of values at any points. However, this greatly increases the difficulty of learning, as it requires learning the entire solution in the spatiotemporal domain. In contrast, discrete PINNs only need to learn the field-to-field mapping. The AD used in continuous PINNs eliminates the error in differentiation, but also increases the computational cost. Additionally, continuity makes it difficult for continuous PINNs to learn large gradients and discontinuous features, which are easier for discrete PINNs. Moreover, discrete PINNs are easier to implement hard constraints on PDEs, which will be elaborated in the part of "soft vs hard constraints" in this section. The choice between continuous PINNs and discrete PINNs should depend on the specific task type. For example, continuous PINNs are more suitable for learning continuous spatiotemporal solutions, reconstructing non-grid-based distributed data, and performing zero-error differentiation. Discrete PINNs are more suitable for tasks such as surrogate modeling of chemical kinetics (Section 4.2), super-resolution (Section 5.5), and AR simulations 119.
@@
Complete/detailed vs reduced/simplified physics. As mentioned in Sections 8.1.1 and 8.1.2, not all studies on PIML for combustion adopt complete/detailed physical models, such as detailed reaction mechanisms, complete PDEs, non-averaged NS equations, and closed turbulence terms. The choice between complete/detailed or reduced/simplified physical models should depend on the specific circumstances. For forward problems, the physical model can be simplified but must be complete, otherwise the problem will be ill-posed. For inverse problems with data, the choice is more flexible, and either a complete/detailed or a reduced/simplified physical model can be used. In fact, sometimes using only a few physical constraints can greatly enhance the overall performance. For example, some studies in Sections 4.2 and 4.3.1 show that using only species and element conservations can achieve a significant improvement. When there are discrepancies between the physics, model, and data, a reduced/simplified physical model can be used. See Section 8.1.2 for details. When both complete models and data are available, comparative studies can be conducted to identify which models have minimal impact on the results and can be omitted, thereby reducing computational costs, as demonstrated in Reference 313.
@@ 8.3 Opportunities for combustion and AI
  • Integration of advanced AI methods. As introduced in this paper, various AI methods haven been used in PIML for combustion, including the basic MLP, KAN 283, NO, NODE, GAN, LSTM, AE, self-supervised learning, and transfer learning 304. However, there are still many advanced AI methods that have potential for application, such as DM, reinforcement learning, foundation models, and PDE discovery. Among these, DM has already been applied to fluid mechanics 169, 444, 445 and combustion 304, 446. Foundation models are a hot topic due to their notable applications in the NLP field, including LLMs such as GPT 25, Llama 447, and DeepSeek 448. In the SciML field, foundation models are a current key direction 89, 449. Some foundation models for PDEs have been developed, including the PDEformer 450, 451, UPS 452, Poseidon 453, PROSE-PDE 454, and PROSE-FD for fluid PDEs 455. Therefore, establishing foundation models for combustion is a feasible and promising direction 46, which also needs to be informed by combustion physics.
@@
  • Combustion-inspired AI. Physics-inspired AI paradigms are an important topic, in which AI methods are often proposed or improved by drawing inspiration from physical laws, phenomena, or principles. The inspiration can be achieved in many ways, including embedding physical laws, mimicking physical processes, and borrowing physical principles. Therefore, physics-inspired AI includes physics-informed AI, or it can be considered as physics-informed AI in a broad sense. Jiao et al. 456 have reviewed the AI methods that are inspired by classical mechanics, electromagnetism, statistical physics, and quantum mechanics. Additionally, two currently popular generative models are also inspired by physics: DM 163, 164 is inspired by nonequilibrium thermodynamics, while the continuity equation in the flow matching 457 originates from fluid mechanics. Therefore, it is natural to ask: can we develop combustion-inspired AI? In fact, CRNN can be regarded as one such effort 78, with its network architecture designed based on the law of mass action and Arrhenius law. However, given the abundance of combustion laws, phenomena, and principles, there are still many possibilities to explore, which may give rise to some revolutionary AI methods.
@@ Acknowledgments
This work is supported by the National Key R&D Program of China (No. 2021YFA0716202), and the National Natural Science Foundation of China (No. 62273149).
@@ Appendix A A code demo of PIML for combustion
[…]
@@
[…]
@@
[…]
@@
[…]
A0 A code demo for the PDEs of 2D unsteady combustion reacting flows using the DeepXDE, with PyTorch as the backend. The Dufour and Soret effects are neglected.

### v3

@@ 8.1.1 Limitations reinforced by combustion
To deal with these difficulties, the methods introduced in Section 2.3 should be appropriately used and improved. For example, the Fourier multiscale embedding can be used to enhance the multiscale representation 130, which has already been adopted in PIML for combustion 304. However, the introduction of hyperparameters is another minor issue, which can be resolved if the scale information is known. Moreover, there is no guarantee that this method will be effective on problems with a wider range of scales. In such cases, it is necessary to combine the domain decomposition methods, which are particularly important for large-domain and long-time problems. Domain decomposition can also facilitate parallel computation, thus accelerating the simulation. It is recommended to use knowledge-based domain decomposition if the approximate distribution characteristics of the solution are known in advance. Examples include combustors with reactant, mixing, reaction, and product regions. Adaptive domain decomposition can also be proposed for more complex problems. Unlike common domain decomposition methods, for combustion problems, each subdomain can use not only different networks but also different physical laws, which is similar to the adaptive combustion models 50. To deal with the multiphysics issue, imposing hard constraints on certain physical laws can always enhance the physical consistency and generalization ability. As for the ill-conditioning issue, some preconditioning strategies should be adopted, such as the normalization. For a deeper understanding, the ill-conditioning analysis should also be conducted for PIML solving combustion PDEs, which can be enlightened by the papers mentioned in Section 2.3.4. Although the reaction stiffness has been successfully solved in reaction ODE problems (see Section 4.1), it has not been thoroughly solved in reacting flow problems. As summarized in Table 5, when using the direct finite-rate chemistry in PIML, the most detailed reaction mechanism is only a two-step mechanism. Other studies either discard, simplify, tabulate, pre-learn, or online-learn the reaction part. Among these, only tabulating and pre-learning can represent the known detailed mechanism and may serve as feasible ways to tackle the stiffness issue in reacting flows, as long as they are differentiable. In fact, tabulating and pre-learning with NNs can be viewed as preconditioning strategies and can be categorized as building surrogate models of chemical kinetics, as mentioned in Section 4.2, not for CFD, but for PIML.
@@ 8.1.2 Discrepancies between physics, (physical) model, and data
As exemplified in Sections 4 and 5, to deal with these discrepancies between physics, model, and data, three strategies can be adopted: (1) Use reduced models, i.e., keep only the certain models and discard the uncertain parts. For example, the mass/continuity equation was adopted in most studies, as shown in Table 5, because it is certain, simple, and do not involve source terms that may need to be modeled for other equations. For the same reason, the energy and species equations are discarded in many studies. (2) Use simplified models. Examples include using RANS or LES models for turbulence and using simplified reaction mechanisms for complex reactions. Filtered PDEs have been proven to be effective for PIML to handle real-world data 207. (3) Learn the models, especially the turbulence and reaction closures, which have been illustrated in References 311, 109, 110, 312. These works demonstrate that the potential of PIML in inverse problems is not limited to field reconstruction and parameter inference, but can also assist in physical modeling. In addition to designing strategies to address these discrepancies, it is also necessary to have methods to measure these discrepancies, i.e., uncertainty quantification (UQ) 216, 47. The methods for quantifying data and model uncertainties mentioned in Section 2.3.4 can be applied to PIML for combustion.
@@ 8.1.3 Reproducibility
Reproducibility is the foundation of academic research, ensuring the reliability, transparency, and implementability of the adopted methods in the papers, thereby guaranteeing the credibility of their conclusions. Many fields face reproducibility issues, or the so-called reproducibility crisis 441, 442, and we believe that PIML for combustion is no exception. According to Reference 443, overoptimism exists in ML for fluid-related PDEs, which is caused by weak baselines and reporting biases. The former does not apply to most studies on PIML for combustion because their goal is not to exceed a baseline model. However, reporting bias is highly likely to exist, which includes selective reporting, outcome switching, and publication bias 443. In addition to the reporting bias, the reproducibility issues can also be caused by other pitfalls, including the sensitivity to hyperparameters, ambiguity in implementation details, and inaccessibility of data and codes. As demonstrated in Tables 4 and 5, most studies on PIML for combustion do not share codes, which is very different from the open source ecosystem of the AI community.
To deal with the reproducibility issues, McGreivy et al. 443 call for bottom-up cultural changes to minimize biased reporting and top-down structural reforms to reduce perverse incentives. We briefly outline some solutions at the researcher level: (1) Provide detailed model configurations, parameter settings, and evaluation metrics. (2) Improve data accessibility by either making data available or using publicly available data, such as those in Table 6. When establishing public datasets, it is preferable to follow the principles of FAIR (Findability, Accessibility, Interoperability, and Reusability) 444. When establishing public benchmarks, some suggestions can be found in Reference 356. (3) It is preferable to provide the codes, as we have done in References 115, 319 and Listing 1. When developing public software, some suggestions can be found in Reference 445. (4) Report not only positive results but also negative results. (5) Conduct replicate experiments and report the mean and variance of the results to demonstrate that the impact of randomness is negligible. (6) Conduct hyperparameter sensitivity analysis. For hyperparameters that are highly sensitive, sufficient explanation should be provided. Finally, we emphasize that the reproducibility issues can also be alleviated by the establishment of an open, active, and rigorous community for PIML for combustion, which involves the standardization of processes, tools, and definitions, as well as the open sourcing of data, codes, and tools. There is still a long way to go in this regard.
@@ 8.2 Discussions on implementation choices
Forward vs inverse problems. As shown in Tables 4 and 5, PIML is currently used more for inverse problems than for forward problems, and this is not limited to the field of combustion. There are many reasons for this: (1) Inverse problems are usually simpler than forward problems when both types of problems can be studied because of the involvement of data. When solving forward problems fails, inverse problems are the only option. (2) Inverse problems are more common than forward problems in practice. Forward problems have strict requirements and need complete equations and IC/BCs, while inverse problems have more relaxed requirements. Therefore, it is not that researchers do not study forward problems, but that the specific problems themselves are inverse problems. (3) Inverse problems have richer application significance than forward problems in terms of field reconstruction, parameter inference, physical modeling, and kinetics discovery. Forward problems often only involve simulation, which can also be achieved by traditional numerical methods. (4) Inverse problems are better suited to leveraging the advantages of PIML than forward problems. It is well known that the overall performance of PIML in forward problems is far inferior to that of traditional numerical methods 443. However, one advantage of PIML is that it can easily couple data and incomplete equations to solve inverse problems, which is not straightforward for traditional numerical methods. In fact, the data-driven characteristic is an inherent advantage of ML, which can be leveraged in PIML. Nevertheless, the forward problems still have two important aspects: (1) In terms of academic research, the forward problem is the basis of the inverse problem. Research on the forward problems helps deepen our understanding of PIML, which in turn helps us understand the inverse problems. Furthermore, many proposed methods on the forward problems in terms of representation models, optimization processes, and PDE systems can be seamlessly transferred to the inverse problems. (2) In terms of practical applications, although PIML is inferior to traditional solvers in solving single PDE, it has demonstrated advantages in parametric/surrogate solving. Examples can be found in Section 2.3.4, Section 2.4, and Reference 304. Compared with traditional solvers, it is more efficient, provides a continuous representation in the parameter space, and can easily perform parameter sensitivity analysis. Note that parametric/surrogate modeling can also be implemented in the way of inverse problems, where the data can be utilized. To sum up, both types of problems are important. For researchers and engineers, the choice between forward problems and inverse problems should depend on the specific circumstances and objectives.
@@
Continuous vs discrete PINNs. As defined in Section 2.3.1, continuous PINNs follow the coordinate-to-variable mapping way, i.e., $\widehat{\mathbf{u}}=\mathrm{NN}(\mathbf{z})$, while discrete PINNs follow the field-to-field mapping way, i.e., $\widehat{\mathbf{U}}_{\mathrm{out}}=\mathrm{NN}(\mathbf{U}_{\mathrm{in}})$. Both have their own advantages and disadvantages 139. Continuous PINNs provide a continuous representation of the solution, making them mesh-independent and allowing for easy query of values at any points. However, this greatly increases the difficulty of learning, as it requires learning the entire solution in the spatiotemporal domain. In contrast, discrete PINNs only need to learn the field-to-field mapping. The AD used in continuous PINNs eliminates the error in differentiation, but also increases the computational cost. Additionally, continuity makes it difficult for continuous PINNs to learn large gradients and discontinuous features, which are easier for discrete PINNs. Moreover, discrete PINNs are easier to implement hard constraints on PDEs, which will be elaborated in the part of "soft vs hard constraints" in this section. The choice between continuous PINNs and discrete PINNs should depend on the specific task type. For example, continuous PINNs are more suitable for learning continuous spatiotemporal solutions, reconstructing non-grid-based distributed data, and performing zero-error differentiation. Discrete PINNs are more suitable for tasks such as surrogate modeling of chemical kinetics (Section 4.2), SR (Section 5.5), and AR simulations 119.
@@
Complete/detailed vs reduced/simplified physics. As mentioned in Sections 8.1.1 and 8.1.2, not all studies on PIML for combustion adopt complete/detailed physical models, such as detailed reaction mechanisms, complete PDEs, non-averaged NS equations, and closed turbulence terms. The choice between complete/detailed or reduced/simplified physical models should depend on the specific circumstances. For forward problems, the physical model can be simplified but must be complete, otherwise the problem will be ill-posed. For inverse problems with data, the choice is more flexible, and either a complete/detailed or a reduced/simplified physical model can be used. In fact, sometimes using only a few physical constraints can greatly enhance the overall performance. For example, some studies in Sections 4.2 and 4.3.1 show that using only species and element conservations can achieve a significant improvement. When there are discrepancies between the physics, model, and data, a reduced/simplified physical model can be used. See Section 8.1.2 for details. When both complete models and data are available, comparative studies can be conducted to identify which models have minimal impact on the results and can be omitted, thereby reducing computational costs, as demonstrated in Reference 312.
@@ 8.3 Opportunities for combustion and AI
  • Integration of advanced AI methods. As introduced in this paper, various AI methods haven been used in PIML for combustion, including the basic MLP, KAN 283, NO, NODE, GAN, LSTM, AE, self-supervised learning, and transfer learning 308. However, there are still many advanced AI methods that have potential for application, such as DM, reinforcement learning, foundation models, and PDE discovery. Among these, DM has already been applied to fluid mechanics 169, 446, 447 and combustion 308, 448. Foundation models are a hot topic due to their notable applications in the NLP field, including LLMs such as GPT 25, Llama 449, and DeepSeek 450. In the SciML field, foundation models are a current key direction 89, 451. Some foundation models for PDEs have been developed, including the PDEformer 452, 453, UPS 454, Poseidon 455, PROSE-PDE 456, and PROSE-FD for fluid PDEs 457. Therefore, establishing foundation models for combustion is a feasible and promising direction 46, which also needs to be informed by combustion physics.
@@
  • Combustion-inspired AI. Physics-inspired AI paradigms are an important topic, in which AI methods are often proposed or improved by drawing inspiration from physical laws, phenomena, or principles. The inspiration can be achieved in many ways, including embedding physical laws, mimicking physical processes, and borrowing physical principles. Therefore, physics-inspired AI includes physics-informed AI, or it can be considered as physics-informed AI in a broad sense. Jiao et al. 458 have reviewed the AI methods that are inspired by classical mechanics, electromagnetism, statistical physics, and quantum mechanics. Additionally, two currently popular generative models are also inspired by physics: DM 163, 164 is inspired by nonequilibrium thermodynamics, while the continuity equation in the flow matching 459 originates from fluid mechanics. Therefore, it is natural to ask: can we develop combustion-inspired AI? In fact, CRNN can be regarded as one such effort 78, with its network architecture designed based on the law of mass action and Arrhenius law. However, given the abundance of combustion laws, phenomena, and principles, there are still many possibilities to explore, which may give rise to some revolutionary AI methods.
@@ Acknowledgments
This work is supported by the National Key R&D Program of China (No. 2021YFA0716202), the National Natural Science Foundation of China (No. 62273149), the Science and Technology Plan Project of Yunnan Province (No. 202302AQ370003-2), and the Creative Seed Fund of Shanxi Research Institute for Clean Energy, Tsinghua University.
@@ Appendix A A code demo of PIML for combustion
u, v, p, T, Ys = uvpTYs[:, 0:1], uvpTYs[:, 1:2], uvpTYs[:, 2:3], uvpTYs[:, 3:4], uvpTYs[:, 4:]
# compute the derivatives
@@
# infer the correction coefficient for viscosity (cmu) in inverse problems, if users want
@@
# compute some terms in the continuity and momentum equations
@@
return [continuity, momentum_x, momentum_y, energy] + [species[:, k:k+1] for k in range(n_spe)]
Listing 1: A code demo for the PDEs of 2D unsteady combustion reacting flows using the DeepXDE, with PyTorch as the backend. The Dufour and Soret effects are neglected.